OpenStudy (anonymous):

What is 0.5(x-3)+(1.5-x)=5x

terenzreignz (terenzreignz):

Simplify what needs to be simplified. That begins by getting rid of the parentheses.

terenzreignz (terenzreignz):

(correctly, of course)

OpenStudy (anonymous):

So I distribute 0.5 first?

terenzreignz (terenzreignz):

Yes. And as for the +(1.5 - x) bit, that's like distributing 1, which just means, get rid of the parentheses :)

OpenStudy (anonymous):

So you will get 0.5x-1.5+1.5-x+5x

terenzreignz (terenzreignz):

You mean = 5x at the last bit, right? typo?

OpenStudy (anonymous):

Yes = 5x .. huge typo

terenzreignz (terenzreignz):

0.5x-1.5+1.5-x = 5x Okay, can you take it from here?

OpenStudy (anonymous):

I can but what is 0.5x-x ..

terenzreignz (terenzreignz):

Ahh... I see where this might be confusing... you want to multiply both sides by 10?

terenzreignz (terenzreignz):

It doesn't change the answer, but it might make it easier for you...

OpenStudy (anonymous):

Both sides of the equal side?

terenzreignz (terenzreignz):

Yes, of course. It's only fair :) Multiplication property of equality...

OpenStudy (anonymous):

So 0.5x times 10 will become 5x then subtract by x it will equal 4x?

terenzreignz (terenzreignz):

Whoops... you also have to multiply the x by 10. You multiply EVERYTHING by 10 so that the equality remains.

OpenStudy (anonymous):

Aww okay .. So how would I write the whole equation out?

terenzreignz (terenzreignz):

\[\Large 0.5x-1.5+1.5-x = 5x\] \[\Large \color{red}{10}(0.5x-1.5+1.5-x) =\color{red}{10}( 5x)\] We just multiplied 10 on both sides... not really necessary, but it makes it easier to solve :)

terenzreignz (terenzreignz):

So, distribute on the left, multiply on the right, what do you get?

OpenStudy (anonymous):

5x-15+15-10x=50x

terenzreignz (terenzreignz):

Yeah, does that allay your confusion? :)

OpenStudy (anonymous):

Yes .. A lot!!

terenzreignz (terenzreignz):

Okay, try solving it now :)

OpenStudy (anonymous):

When I get down to -5x=50x I just divide?

terenzreignz (terenzreignz):

You can't, because they both have x. You have to bring all numbers with x to one side.

terenzreignz (terenzreignz):

err... all *terms* with x.

OpenStudy (anonymous):

Soo how again do I that.

terenzreignz (terenzreignz):

Well, Let's bring the -5x to the right side... it's a minus, so what do we do to get rid of it?

OpenStudy (anonymous):

Add it on both sides

terenzreignz (terenzreignz):

Add...what?

OpenStudy (anonymous):

-5x

terenzreignz (terenzreignz):

No, that gets you nowhere... -5x -5x = 50x - 5x -10x = 45x

OpenStudy (anonymous):

Okay that solves everything .. I'm finally good now

terenzreignz (terenzreignz):

okay... if you're sure... what's your answer?

OpenStudy (anonymous):

Should I get a zero for 0=4.5x .. this is what my teacher got last period.

terenzreignz (terenzreignz):

Yes! That is correct ^_^

OpenStudy (anonymous):

Omg .. Yes!! .. I feel so smart now lol!

terenzreignz (terenzreignz):

I'm sure you are :D

OpenStudy (anonymous):

Thank you!